# custom.bib

### v1

@@ C.2 Evaluation metrics
This section explains the evaluation metrics used in the extended ZsRE[Yao, knowedit] and $\text{Wiki}_{counterfact}$ [ripple-effects, knowedit]. Generally, they adopt four metrics: reliability, generality, portability, and locality. Given an initial base model $f_{\theta}$, a post-edit model $f_{\theta^{\prime}}$, and a set of edit instances $(x_{t},y_{t})\in\{(x_{t},y_{t})\}$, the reliability is computed as the average accuracy of the edit cases:
@@
Despite the editing, those instances that are irrelevant to the edit cases $(\hat{x_{t}},\hat{y_{t}})\in\{O(x_{t},y_{t})\}$ should not be affected. This evaluation is called locality (also known as specificity) and is measured by the proportion of unchanged predictions between the initial model and the post-edit model:
$\mathbb{E}_{(\hat{x_{t}},\hat{y_{t}})\in\{O(x_{t},y_{t})\}}\{f_{\theta^{\prime}}(\hat{x_{t}})=f_{\theta}(\hat{x_{t}})\}\ .$ (12)
For the extended ZsRE [Yao, knowedit] and $\text{Wiki}_{counterfact}$ [ripple-effects, knowedit], we follow the setting in the original paper and combine reliability and generality to the Edit Success rate.
@@ C.3 Baseline implementation details
Unless otherwise specified, the baselines are implemented by using the EasyEdit framework [EasyEdit].
@@
We follow the procedure implemented in previous work [rome, memit, Yao, knowedit] to fine-tune a specific layer from the model. We select layer 13 for Llama-3.2-1B and layer 27 for Qwen2.5-7B. For both models, we adopt the learning rate of $5e^{-4}$ and the number of optimization steps 25.
@@
For both models, we use LoRA [LoRA] to update the query and key projection matrix of the models, with rank set to 8, $\alpha$ set to 32, the dropout rate 0.1, and the learning rate $5e^{-3}$. The number of updating steps is set to 70 for Llama-3.2-1B and 60 for Qwen2.5-7B.
@@
ROME [rome] treats the FFN part of the LLMs as a key-value association and updates a pre-located layer by directly inserting an optimized key-value pair. We update the layer 5 for both Llama-3.2-1B and Qwen2.5-7B, and adopt 25 optimization steps for Llama-3.2-1B and 20 optimization steps for Qwen2.5-7B, with both learning rate $5e^{-5}$.
@@
R-ROME [r-rome] is another version of ROME [rome] with modified code implementation. We use the same hyperparameters as ROME.
@@
KN [kn] hypothesize that factual knowledge is stored in FFN memories and expressed by knowledge neurons. For both models, we use the threshold of 0.2 for knowledge attribution scores and 0.4 for the threshold of the prompts sharing percentage.
@@
GRACE [grace] adopts a discrete codebook to memorize the edits as key-value pairs. We set the location of the codebook layer 13 and 18 for Llama-3.2-1B and Qwen2.5-7B, respectively. Surprisingly, the $\epsilon$ value used in the original paper (1-3) seems insufficient for the complex editing experiments in this paper. Therefore, we increase it to 50. The number of optimization steps for the value vector is set to 100.
@@
IKE [ike] maintains an explicit memory for edits and retrieves them via K-nearest neighbors. The retrieved edits are then used to construct demonstrations, which are then prefixed to the input to edit the behavior. In the experiments, we set $K=16$.
@@
MEND [mend] trains an additional meta-network to predict a new rank-one update to the input gradient. In this paper, we train each model using ZsRE [zsre] and COUNTERFACT [memit], and adopt the ZsRE-trained model to MQuAKE and the extended ZsRE and COUNTERFACT-trained model for $\text{Wiki}_{counterfact}$.
@@
SERAC [serac] employs an explicit edit-instance memory, an additional trained scope classifier, and a trained counterfactual model. The scope classifier is responsible for determining whether an input is relevant to the edits in the memory. The input is fed to the counterfactual model once the input is deemed as relevant to memorized edits and the original model otherwise. We use the distilbert-base-cased [distil-bert] model as the scope classifier, and train it using the training set of ZsRE [zsre] and COUNTERFACT [memit] from EasyEdit [EasyEdit]. Following [dune], we use instruction-tuned models (Llama-3.2-1B-Instruct [^11] https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct and Qwen2.5-0.5B [^12] https://huggingface.co/Qwen/Qwen2.5-0.5B-Instruct) for the counterfactual model.
@@
MEMIT [memit] is the extension of ROME [rome] that supports a batch of edits at a time. Unlike ROME, which only updates a single pre-located layer, MEMIT spreads the update to a set of identified layers. We apply changes to layers 4-8 for both models. Following the settings in the original paper, we set $\lambda$, the hyperparameter that balances the weighting of new and old associations, to $1.5\times 10^{4}$.
@@
EMMET [emmet] is a batched-version of ROME [rome]. Similar to ROME, we edit the layer 5 for both models, and set $\lambda=1e^{5}$.
Appendix D Societal impacts

### v2

@@ C.2 Evaluation metrics
This section explains the evaluation metrics used in the extended ZsRE[43, 46] and $\text{Wiki}_{counterfact}$ [2, 46]. Generally, they adopt four metrics: reliability, generality, portability, and locality. Given an initial base model $f_{\theta}$, a post-edit model $f_{\theta^{\prime}}$, and a set of edit instances $(x_{t},y_{t})\in\{(x_{t},y_{t})\}$, the reliability is computed as the average accuracy of the edit cases:
@@
Despite the editing, those instances that are irrelevant to the edit cases $(\hat{x_{t}},\hat{y_{t}})\in\{O(x_{t},y_{t}),f_{\theta}(x_{t})=y_{t}\}$ should not be affected. This evaluation is called locality (also known as specificity) and is measured by the proportion of unchanged predictions between the initial model and the post-edit model:
[…]
For the extended ZsRE [43, 46] and $\text{Wiki}_{counterfact}$ [2, 46], we follow the setting in the original paper and combine reliability and generality to the Edit Success rate.
@@ C.3 Baseline implementation details
Unless otherwise specified, the baselines are implemented by using the EasyEdit framework [39].
@@
We follow the procedure implemented in previous work [24, 25, 43, 46] to fine-tune a specific layer from the model. We select layer 13 for Llama-3.2-1B and layer 27 for Qwen2.5-7B. For both models, we adopt the learning rate of $5e^{-4}$ and the number of optimization steps 25.
@@
For both models, we use LoRA [16] to update the query and key projection matrix of the models, with rank set to 8, $\alpha$ set to 32, the dropout rate 0.1, and the learning rate $5e^{-3}$. The number of updating steps is set to 70 for Llama-3.2-1B and 60 for Qwen2.5-7B.
@@
ROME [24] treats the FFN part of the LLMs as a key-value association and updates a pre-located layer by directly inserting an optimized key-value pair. We update the layer 5 for both Llama-3.2-1B and Qwen2.5-7B, and adopt 25 optimization steps for Llama-3.2-1B and 20 optimization steps for Qwen2.5-7B, with both learning rate $5e^{-5}$.
@@
R-ROME [10] is another version of ROME [24] with modified code implementation. We use the same hyperparameters as ROME.
@@
KN [3] hypothesize that factual knowledge is stored in FFN memories and expressed by knowledge neurons. For both models, we use the threshold of 0.2 for knowledge attribution scores and 0.4 for the threshold of the prompts sharing percentage.
@@
GRACE [12] adopts a discrete codebook to memorize the edits as key-value pairs. We set the location of the codebook layer 13 and 18 for Llama-3.2-1B and Qwen2.5-7B, respectively. Surprisingly, the $\epsilon$ value used in the original paper (1-3) seems insufficient for the complex editing experiments in this paper. Therefore, we increase it to 50. The number of optimization steps for the value vector is set to 100.
@@
IKE [48] maintains an explicit memory for edits and retrieves them via K-nearest neighbors. The retrieved edits are then used to construct demonstrations, which are then prefixed to the input to edit the behavior. In the experiments, we set $K=16$.
@@
MEND [28] trains an additional meta-network to predict a new rank-one update to the input gradient. In this paper, we train each model using ZsRE [20] and COUNTERFACT [25], and adopt the ZsRE-trained model to MQuAKE and the extended ZsRE and COUNTERFACT-trained model for $\text{Wiki}_{counterfact}$.
@@
SERAC [29] employs an explicit edit-instance memory, an additional trained scope classifier, and a trained counterfactual model. The scope classifier is responsible for determining whether an input is relevant to the edits in the memory. The input is fed to the counterfactual model once the input is deemed as relevant to memorized edits and the original model otherwise. We use the distilbert-base-cased [38] model as the scope classifier, and train it using the training set of ZsRE [20] and COUNTERFACT [25] from EasyEdit [39]. Following [1], we use instruction-tuned models (Llama-3.2-1B-Instruct [^10] https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct and Qwen2.5-0.5B [^11] https://huggingface.co/Qwen/Qwen2.5-0.5B-Instruct) for the counterfactual model.
@@
MEMIT [25] is the extension of ROME [24] that supports a batch of edits at a time. Unlike ROME, which only updates a single pre-located layer, MEMIT spreads the update to a set of identified layers. We apply changes to layers 4-8 for both models. Following the settings in the original paper, we set $\lambda$, the hyperparameter that balances the weighting of new and old associations, to $1.5\times 10^{4}$.
@@
EMMET [11] is an unification of ROME [24] and MEMIT [25]. Similar to ROME, we edit the layer 5 for both models, and set $\lambda=1e^{5}$.
RAG
We adopt bge-base-en-v1.5 [41] as our retriever. For batch editing, we treat the corresponding batch number of edits as our corpus, and retrieve the 10 most relevant edits for each testing query.
InComeS
We apply cross-attention operations only for the second half of the model’s layers since we found that the gist KV cache from the first half is not informative enough to allow effective edit selection. For the calculation of cross-attention during inference, we adopt a temperature of $T=0.45$ to the logits before softmax, which is found to be helpful for effective editing.
Appendix D Further Analysis
D.1 RAG vs. InComeS
Problem settings
InComeS and RAG target different problems. In RAG, the query is given and used to retrieve the relevant documents via a retrieval model before decoding. However, InComeS does not make such an assumption, and it "edits" the model with the provided knowledge before seeing any actual queries.
Methodology
From the perspective of methodology, InComeS conducts selection mechanisms in a way that is better integrated into the LM and with a finer granularity compared to RAG methods. First, RAG inputs the full query for retrieval to select relevant documents, and this process needs an extra retrieval model; in contrast, our method requires no extra models or retrieval steps and directly integrates the selection mechanism into our base LM. Moreover, our method dynamically performs selection for each token individually, which has a much finer granularity than the query-based selection in RAG. This supports a wider range of applications.
Experiments on Multi-hop edits
Despite the differences between our method and RAG, we compare our method with RAG to demonstrate our method’s superiority over the complex editing scenarios (Table 8). InComeS outperforms RAG in almost all cases for both single editing and batch editing. Note that the result of RAG is the same as the result of ICL in single editing, which does not need retrieval.
Method Model Single Editing Batch Editing 2-edits 3-edits 4-edits 2-edits 3-edits 4-edits Base Llama-3.2-1B 41.79 43.51 31.58 41.79 43.51 31.58 RAG 59.23 59.00 51.63 40.65 46.78 43.58 ICL 59.23 59.00 51.63 50.06 49.86 42.37 InComeS 71.19 72.17 72.62 53.93 52.79 52.73 Base Qwen2.5-7B 44.08 44.14 30.62 44.08 44.14 30.62 RAG 69.76 76.91 74.54 42.91 46.37 46.15 ICL 69.76 76.91 74.54 53.53 50.54 44.77 InComeS 66.46 71.24 76.54 55.13 53.48 47.91
Table 8: Results for RAG on MQuAKE [49].
D.2 Side effect
We present a side effect analysis of our method in this section (Table 9). We test the editing side effect under three different numbers of edits (0, 0.1k, 1k) on MMLU [13] benchmark, which consists of 57 tasks across 4 domains, namely Social science, Humanities, STEM, and others. The results indicate that increasing the number of edits does not significantly harm the model’s general capability (lines 3 - 5 and 7 - 9 in Table 9), demonstrating the potential scalability of our method. The continuous pre-training brings an inevitable modest side effect to the model (lines 2 - 3 and 6- 7 in Table 9).
Method Model Social Sciences Humanities STEM Other Average Base Llama-3.2-1B 25.09 27.31 25.87 27.49 26.44 InComeS - w/ no edits 23.18 25.91 22.66 26.06 24.45 InComeS - w/ 0.1k edits 22.45 25.01 22.66 24.62 23.68 InComeS - w/ 1k edits 22.91 25.30 22.73 24.11 23.76 Base Qwen2.5-7B 72.72 69.19 63.42 69.72 68.76 InComeS - w/ no edits 68.06 65.68 59.58 66.39 64.93 InComeS - w/ 0.1k edits 64.64 60.65 56.84 63.39 61.38 InComeS - w/ 1k edits 63.86 62.08 57.58 62.88 61.60
Table 9: Side effect evaluation on MMLU [13].
D.3 Efficiency analysis
Compared to many traditional editing methods that require model backward calculation, our method only requires one single forward pass for each editing context. In comparison to ICL, which needs to encode the entire concatenated edit context, our approach enables parallel encoding of multiple edits, leading to great efficiency gains for the encoding (prefilling) stage. In addition, the compressed context also accelerates the decoding phase compared to the ICL decoding with prefilled KV cache. Here, we provide an analysis of our method’s efficiency advantage over traditional ICL for both the encoding and decoding stages.
Encoding
Assume we have $N$ edits and each edit has a Length of $L$. For ICL prefilling, it has to encode the whole sequence with length $N\times L$. However, for InComeS, each edit is processed individually, and it encodes edits in parallel. In this case, it encodes a batch of $N$ edits with length $L$. Thanks for the highly optimized GPU parallel computation, such a feature approximately reduces the time consumption by $N$ times.
Decoding
Suppose we have $N$ compressed gists, which corresponds to $N$ individual edits with length $L$ and $N\times L$ tokens whose KV caches have been prefilled. For each decoding position, the ICL self-attention needs to compute a matrix with size $1\times N\times L$. However, InComeS only needs to calculate the gist cross-attention matrix with size $1\times N$. This roughly accelerates the decoding by $L$ times.
Appendix E Societal impacts